OpenStudy (anonymous):

EVALUATE THE EXPRESSION let d = 2 (sorry for caps) 3+2d over 14

OpenStudy (anonymous):

ty!

OpenStudy (anonymous):

Is it (3+2d)/14 OR: 3 + (2d/14)?

OpenStudy (anonymous):

3 + (2d/14)

OpenStudy (anonymous):

oh okay. 3 + (4/14) = 3 + 2/7 3(7)/7 + 2/7 = 21/7 + 2/7 = 23/7

OpenStudy (anonymous):

I GET IT! ty!

OpenStudy (anonymous):

I'm just pointing out that the order of operation makes a BIG difference :-)

OpenStudy (anonymous):

it doess haha :)))))